# Building a continuous benchmarking ecosystem in bioinformatics

### v2

@@ Introduction
Benchmarking (defined below) is a critical step in method development. Such studies are typically published as either ‘methods-development papers’ (MDPs) where new methods are compared against existing ones; or ‘benchmark-only papers’ (BOPs), where a set of existing methods are compared in a more neutral way, as defined in [1]. A benchmarking study can be run by a single person or a small group, or it can be organized as a ‘challenge’ or within a hackathon. In this work, we present a perspective on building a robust benchmarking ecosystem in bioinformatics, coming primarily from the methodologist’s viewpoint (researchers developing computational tools). While our focus is bioinformatics, we aim to abstract principles and connect to similar challenges in adjacent fields.
In our discussion, we outline concepts and specifications for an ideal benchmarking system; we examine philosophical issues like gatekeeping and trust; and provide some technical recommendations, such as standardization.
@@
We define a benchmark as a conceptual framework to evaluate the performance of computational methods for a given task, though it can extend to assessing other aspects, like data-generating technologies. A benchmark thus requires a well-defined task and a definition of correctness, or ground-truth, to be precisely defined in advance. A benchmark then consists of benchmark ‘components’, such as simulations to generate datasets, preprocessing steps, methods, and metrics.
@@
Can benchmarking be systematized? Benchmarks should be orchestrated and run by benchmarking systems to create ‘artifacts’ (code snapshots, file outputs to be shared, performance outputs). The various elements of code, especially the details on how datasets are preprocessed and how methods are run, should be explicitly available for scrutiny. Such a system should facilitate components to be public, open and allow contributions (e.g., corrections or new components), such that the benchmarks of the future can run more continuously. At the outset of a benchmark, a benchmark ‘definition’ could be envisioned, which could give a formal specification of the entire set of components (and the pattern of artifacts to be generated). The benchmark definition could even be expressed as a single configuration file that specifies the scope and topology of components to include, details of the repositories (code implementations with versions), the instructions to create software environments (accessible across compute architectures), the parameters used, and what components to snapshot for a release. A benchmark definition in terms of layers and their corresponding challenges and opportunities within each layer is illustrated in Figure 1.
@@
Figure 1: Benchmarking conceptual layers and their challenges. The process of benchmarking is multilayered, with layer-specific factors belonging to different categories (colored boxes) and challenges (in red). Hardware: infrastructure and its costs. Data: dataset archival, openness, interoperability, archival, and selection. Software: method implementations, reproducibility, workflow execution, continuous integration and delivery (CI/CD), versioning, quality assurance (QA). Community: Standardization and impartiality (governance, transparency, building trust, long term maintainability). Knowledge: research and meta-research, academic publications.
@@
Benchmarks can also be of high value to scientific journals and funding agencies. Well-executed benchmark studies can be highly utilized and influential, and may point to gaps in the current body of methods, guiding future methodological developments; at the same time, in fast-moving fields like bioinformatics, benchmarks have the tendency to rapidly become stale. Journals and funding bodies also have a vested interest in ensuring that published or funded method developments are performed to a high standard, that unnecessary redundancy is reduced, and that results are FAIR (findable, accessible, interoperable and reusable) for maximal benefit to the community [7]. Thus, quality assurance, neutrality and transparency of benchmarking systems become highly relevant.
@@
Benchmarking as a process deals with tasks other than workflow definition and execution, including: collecting and keeping track of contributions from users; provisioning appropriate hardware; handling a reproducible and efficient software stack; managing storage and access-control lists; rendering results dashboards and summarizing results; versioning code, workflow runs, and files; and providing extensive documentation from defining scope or contribution guidelines, to transparent logging and crediting contributions and disseminating results.
Currently, most benchmarking systems aim to enhance the workflow formalization and execution with strategies to fulfill benchmarking-specific needs. To name a few of these systems, ncbench bundles workflow specification and software management with snakemake to benchmark result visualization with Datavzrd [12]. OpenEBench runs workflows in any workflow language, frequently nextflow, and makes use of openVRE as GUI and means of visualization [13]. OpenProblems uses Viash and nextflow to handle software environments and workflows, respectively, and reports updated leaderboards with custom code [14]. To our knowledge, no benchmarking platform makes use of a fully expressive benchmark definition formalization language covering workflow generation and benchmarking-specific tasks. Such a language would go a long way toward facilitating benchmark conceptualization, community gathering and result sharing.
@@
Traditionally, benchmarks (whether BOP or MDP) are conducted in a decentralized fashion and typically by a small number of researchers (e.g., for MDP, the new method developer). This brings some disadvantages. For example, decentralization can be hugely inefficient; researchers studying the same task would collect reference datasets and code snippets to run methods, and implement a bespoke strategy (e.g., shell scripts) to orchestrate the methods run on datasets. Working together would mean sharing not only datasets and code snippets, but also execution environments (software but perhaps also hardware), alignment on parameters to evaluate and eventually sharing, consolidating, and interpreting results. However, because there are no standards of how benchmarks are implemented [2], another disadvantage of decentralized benchmarks is that, while these components may eventually be openly shared as a product of a research study, the contents of a benchmark are non-interoperable, even across benchmarks of the same task. Furthermore, decentralized benchmarks have been criticized on the basis of ‘inflating model performance’ due to decisions made in preprocessing or parameter tuning [14]; such issues could be mitigated by gathering the ‘wisdom of crowds’ that is implicit in running centralized benchmarks.
@@
Computational reproducibility comes from controlling the triad of data, code, and environment [38]. During benchmarking, a workflow runs code transforming inputs into outputs; execution is affected by the base system (OS, compiler toolchains, libraries) and its configuration. The benchmark definition should control the execution environment.
@@
    Software dependencies cause the biggest impact on replicability and maintainability; archiving all combinations of a large dependency tree, across several OS images, exponentially increases retention and maintenance costs. Making the long-term replicability of a benchmark tractable implies adopting a sane software management system. Several solutions exist, from containerization (e.g., apptainer or singularity) [39] and automation of reproducible and efficient software management (EasyBuild [40] or Spack [41]) to distributing software installations such as CernVM-FS [42], or combinations of the two such as computecanada [43] and EESSI [44].
@@
Even a simple, single-person benchmark comparing methods relies on multiple contributions, from dataset creators and method developers to system software and workflow authors. Hence, data, code and operating system dependencies are all accessible during the benchmarking process. For reproducibility, transparency and to build trust [45], this accessibility and freedom should be extended to others to run (re-run), copy (including fork), distribute (e.g., snapshot), study (e.g., read how someone is implementing a method), change (bug fix, contribute) and improve (extend) benchmarks. Such freedoms are granted (or restricted) by software and content licenses, as are copyright and authorship and attribution [46]. Without explicit licenses, even publicly available code is copyrighted by default, and reuse is restricted (to reproduce, distribute, or create derivative works) [46]. A concise summary of license rights and restrictions, and compatibility across commonly-used open-source licenses can be found in Table 1. Benchmark contributions should be licensed and associated to scholarly names including email and ORCID (as the canonical identifier), hence allowing seamless reuse, transparency and attribution.
@@
Similarly, even if all benchmark code were free software, running it could be hindered by (non-free) proprietary interpreters (e.g., MATLAB). Benchmarking systems should be explicit in allowing or disallowing proprietary software, and be aware and track EULAs (end user license agreements). This conundrum extends to the operating system layer and to system dependencies (e.g., compilers). In terms of benchmark re-runnability, fully free software system dependencies arguably offer the highest reusability, and can be handled in a reproducible manner across operating systems and system architectures (e.g., amd64, arm64) [44].
@@ Acknowledgments
Conceptualization: all. Visualization: BC. Original draft preparation: all. Writing, reviewing and editing: all. Supervision: IM, MDR. Funding acquisition: MDR.
Funding
MDR acknowledges funding from the Swiss National Science Foundation (grants 200021_212940 and 310030_204869) as well as support from swissuniversities P5 Phase B funding (project 23-36_14). CS is supported by the Novartis Research Foundation.
Disclosure statement
No potential conflict of interest is reported by the authors

### v3

@@ Introduction
Benchmarking (defined below) is a critical step in method development. Such studies are typically published as either ‘methods-development papers’ (MDPs) where new methods are compared against existing ones; or ‘benchmark-only papers’ (BOPs), where a set of existing methods are compared in a more neutral way, as defined in [1]. A benchmarking study can be run by a single person (solo benchmarking) or a small group, or it can be organized as a ‘challenge’ or within a hackathon. In this work, we present a perspective on building a robust benchmarking ecosystem in bioinformatics, coming primarily from the methodologist’s viewpoint (researchers developing computational tools). While our focus is bioinformatics, we aim to abstract principles and connect to similar challenges in adjacent fields.
In our discussion, we outline concepts related to workflow automation, benchmark formalization and benchmarking-specific tasks beyond the former, including philosophical issues like gatekeeping and trust; and provide some technical recommendations, such as standardization.
@@
Figure 1: Solo and community benchmarks, workflows, and benchmarking-system-enabled (bs) benchmarks all differ in run properties and their consequences in terms of capabilities and features, including whether: inputs, truths, hidden truths, metrics or run results are available for readers; code is FOSS, software environments are managed, and execution automated; there is a clear track of (human) contributors; results are run in comparable hardware; an entrypoint to results or result interpretation are provided. These run-intrinsic properties have consequences in terms of features (i.e., transparency, forkability, availability for metaanalysis) and technical capabilities (i.e. software versioning or ability to run locally without leaving traces - in incognito mode). Green, blue and red ticks depict available, perhaps available and unavailable features, respectively.
@@
We define a benchmark as a conceptual framework to evaluate the performance of computational methods for a given task, though it can extend to assessing other aspects, like data-generating technologies. A benchmark thus requires a well-defined task and typically a definition of correctness, or ground-truth, to be precisely defined in advance (Figure 1). A benchmark then consists of benchmark ‘components’, such as simulations to generate datasets, preprocessing steps, methods, and metrics.
@@
Can benchmarking be systematized? Benchmarking aims to calculate, collect and report performance metrics of methods aiming to solve a task. Such a process normally requires collecting appropriate data and running methods in an automated manner. Some of these tasks can be automated by workflow systems. Others, such as sharing results in an interactive manner and, particularly, result interpretation and making results reusable by others, go beyond what workflows offer. Similarly, collaborative benchmarking and challenges add extra layers of complexity, including collaboration tracking, hiding ground truths, or the ability to release multiple versions of the same benchmark as it develops (Figure 1). Moreover, scientific benchmarking borrows procedures from the scientific method and from scientific publishing, hence having expectations of fairness, reproducibility, transparency, and trust. Benchmarking systems aim to orchestrate the workflow management, benchmarking and community engagement aspects of the process to generate benchmark ‘artifacts’ (code snapshots, file outputs to be shared, performance outputs) in a systematic way and following good practices and standards. Namely, the various elements of code, especially the details on how datasets are preprocessed and how methods are run, should be explicitly available for scrutiny. Such systems should facilitate components to be public, open and allow contributions (e.g., corrections or new components), such that the benchmarks of the future can run more continuously. At the outset of a benchmark, a benchmark ‘definition’ could be envisioned, which could give a formal specification of the entire set of components (and the pattern of artifacts to be generated). The benchmark definition could even be expressed as a single configuration file that specifies the scope and topology of components to include, details of the repositories (code implementations with versions), the instructions to create software environments (accessible across compute architectures), the parameters used, and what components to snapshot for a release. A benchmark definition in terms of layers and the corresponding challenges and opportunities within each layer is illustrated in Figure 2.
@@
Figure 2: Benchmarking conceptual layers and their challenges. The process of benchmarking is multilayered, with layer-specific factors belonging to different categories (colored boxes) and challenges (in red). Hardware: infrastructure and its costs. Data: dataset archival, openness, interoperability, and selection. Software: method implementations, reproducibility, workflow execution, continuous integration and delivery (CI/CD), versioning, quality assurance (QA). Community: Standardization and impartiality (governance, transparency, building trust, long term maintainability). Knowledge: research and meta-research, academic publications.
@@
Benchmarks can also be of high value to scientific journals and funding agencies. Well-executed benchmark studies can be highly utilized and influential, and may point to gaps in the current body of methods, guiding future methodological developments; at the same time, in fast-moving fields like bioinformatics, benchmarks have the tendency to rapidly become stale. Journals and funding bodies also have a vested interest in ensuring that published or funded method developments are performed to a high standard, that unnecessary redundancy is reduced, and that results are FAIR (findable, accessible, interoperable and reusable) for maximal benefit to the community [7]. Recently published benchmarks and challenges have been conceived as or ported to benchmarking systems to ensure these properties [8, 9, 10]. Thus, quality assurance, neutrality and transparency of benchmarking systems become highly relevant.
Collaborative benchmarking efforts enunciate field-specific challenges and provide a venue for communities to address them. For instance, DREAM [11], GIAB [12] and CASP [13] challenges and benchmarks have demonstrated how collaborative benchmarks can help in moving fields forward.
@@
Benchmarking as a process deals with tasks other than workflow definition and execution, including: collecting and keeping track of contributions from users; provisioning appropriate hardware; handling a reproducible and efficient software stack (the advantages of decoupling environment handling from workflow execution are further discussed in the section ‘Reproducible software environments’); managing storage and access-control lists; rendering results dashboards and summarizing results; versioning code, workflow runs, and files; and providing extensive documentation from defining scope or contribution guidelines, to transparent logging and crediting contributions and disseminating results (Figure 1).
Currently, most benchmarking systems aim to enhance the workflow formalization and execution with strategies to fulfill benchmarking-specific needs. To name a few of these systems, ncbench bundles workflow specification and software management with Snakemake to benchmark result visualization with Datavzrd [18]. OpenEBench runs workflows in any workflow language, frequently nextflow, and makes use of openVRE as GUI and means of visualization [19]. OpenProblems uses Viash [20] and nextflow to handle software environments and workflows, respectively, and reports updated leaderboards with custom code [21]. To our knowledge, no benchmarking platform makes use of a fully expressive benchmark definition formalization language covering workflow generation and benchmarking-specific tasks. Orthogonally, some components of the benchmarking process have been formalized: Viash [20] systematizes workflow components; PROV-O represents provenance data [22]; and Research Objects (RO [23], e.g., RO-Crate [24]) specify a system to store data (files) and their metadata. Reusing pre-existing formalizations and extending them to the benchmarking field would facilitate conceptualization, community gathering and result sharing.
@@
Aside of task-tailored scores (i.e., intrinsic or extrinsic cluster validity measures for clustering tasks), collecting metrics to report computing performance (i.e., wallclock, on-disk read-write operations), requirements (e.g., GPUs, minimum memory) and their scalability depending on the input size (e.g., whether runtimes increase linearly, quadratically, etc) is key during benchmarking. To measure these performance metrics accurately, benchmarking systems ideally would provide means to: i) characterize the hardware running the benchmark, so method runtimes are comparable; ii) track performance (CPU or GPU usage, memory) when running each benchmark component; iii) introduce minimal overhead during profiling; iv) collect variability in performance by running each method several times; and, v) go beyond reporting average resource usage and provide them over execution time (e.g., so embarrassingly-parallel algorithms can be differentiated from largely single-cored runs with few bursts of parallel execution).
@@
Traditionally, benchmarks (whether BOP or MDP) are conducted in a decentralized fashion and typically by a small number of researchers (e.g., for MDP, the new method developer). This brings some disadvantages. For example, decentralization can be hugely inefficient; researchers studying the same task would collect reference datasets and code snippets to run methods, and implement a bespoke strategy (e.g., shell scripts) to orchestrate the methods run on datasets. Working together would mean sharing not only datasets and code snippets, but also execution environments (software but perhaps also hardware), alignment on parameters to evaluate and eventually sharing, consolidating, and interpreting results. However, because there are not many standards of how benchmarks are implemented [2], another disadvantage of decentralized benchmarks is that, while these components may eventually be openly shared as a product of a research study, the contents of a benchmark are non-interoperable, even across benchmarks of the same task. Furthermore, decentralized benchmarks have been criticized on the basis of ‘inflating model performance’ due to decisions made in preprocessing or parameter tuning [21]; such issues could be mitigated by gathering the ‘wisdom of crowds’ that is implicit in running centralized benchmarks.
@@
Compute-intensive benchmarks are often run using cloud or high performance computing, further extending the challenge to ensure hardware homogeneity (for comparable performances) and adding complexity to the benchmarking system capabilities to automate node provisioning (or node selection) and to group executions such that costs are minimized. Some workflow managers, such as Snakemake, provide plugins to execute on the cloud or on-premises queuing systems, effectively decoupling workflow specification and execution.
@@
Authorship, e.g., acknowledging contributions, can be captured, reported and credited using third-party platforms such as Apicuron [38] or by bundling standardized contribution metadata such as the Citation File Format (CFF) [39]. In this way, individual contributions to a benchmark can be acknowledged, browsed, and ideally propagated to academic deliverables (e.g., datasets, publications).
@@
Transparency and clear governance guidelines provide an effective means to detect and react to potential misuse, e.g., by explicit rules to validate results. Other ethical considerations include facilitating safe interactions and fair resource allocation and efficiency. Clear code of conduct statements proactively flagging harassment (e.g., by sex, ethnicity or academic seniority, among others) and provide explicit means to ease collaboration. In terms of efficiency, comparable allocation of resources (e.g., computing power, support) to all contributors, while also being aware of $CO_{2}$ emissions, could also be spelled out in the governance model.
Mature communities are key to sustaining and reflecting the existing challenges in a given field. For instance, CASP, a continuous benchmark by nature, has a established community that defines the scope of critical assessment, input data, target definitions, and composite and normalized metrics to equally weight multiple performance scores [45]. At the continuous benchmark level, CASP keeps track of past performances and introduces metrics to evaluate possible artifacts, i.e., whether progress during a CASP round could arise from different target difficulty [46]. To facilitate community contribution, we believe that benchmarking systems can lower the barrier to contributing, hence making the contribution efforts less likely to be biased.
@@
Computational reproducibility comes from controlling the triad of data, code, and environment [52]. During benchmarking, a workflow runs code transforming inputs into outputs; execution is affected by the base system (operating system, compiler toolchains, libraries) and its configuration. The benchmark definition should control the execution environment.
@@
    Software dependencies cause the biggest impact on replicability and maintainability; archiving all combinations of a large dependency tree, across several OS images, exponentially increases retention and maintenance costs. Making the long-term replicability of a benchmark tractable implies adopting a sane software management system.
The responsibility of handling software environments poses an interesting challenge to benchmarking systems. On one hand, some workflow managers provide means to install and activate software and re-run workflows when software specification changes, even if the workflow topology remains the same. For these, software is the workflow manager’s responsibility. On the other hand, some benchmarks might be designed to compare the same workflow topology run with different software backends (i.e., with conda vs compiled) or with different versions of a dependency (e.g., samtools/htslib releases). Benchmarking systems should provide enough flexibility to make all variants possible (i.e., decouple the benchmark topology from the software backend).
Finally, scientific software management handling is known for its nuances, including abundant dependencies with known incompatibilities and complex installation recipes [53]. Scientific software management is handled via traditional (Linux) package managers (RPMs, DEBs), automated and reproducible installs (such as EasyBuild [54] or Spack [55]), central software stacks and environment modules (i.e., Lmod), conda, containers (e.g., apptainer or Singularity; [56]), or read-only network filesystems providing ready-to-use installations (CernVM-FS) [57]. Further strategies are under active development: EESSI [53] and the Digital Research Alliance of Canada [58] are effectively providing multi-purpose, reusable and efficient software installations via CernVM-FS; conda is speeding up their deployment procedures (via py-rattler/pixi) and allowing compilation to Web assembly (to execute code within a web browser); and apptainer (former Singularity) provides rootless containerization suitable for running in scientific clusters.
Hardware
Software, both from benchmarking systems and from the benchmarking workflow itself (the benchmarked methods), can be run on different hardware, from commodity computers to dedicated hardware, often large servers or academic and commercial clouds. Performance profiling (e.g., how fast a benchmarked method runs) requires access to fair and comparable resources.
Comparable resources can be provisioned on demand in cloud computing, or can be ensured by deploying benchmarking systems in controlled and dedicated hardware. Ideally, performance results (e.g., wallclock time) could be bundled to metadata describing the used hardware (e.g., CPU characteristics).
It is worth noting that CPU microarchitecture influences performance, particularly when using optimized compilers to improve performance of benchmarked methods at the expense of making binaries microarchitecture-specific. This feature is not solved by containerization. Benchmarking systems can automate software installations with equivalent degrees of optimization across benchmarked methods, and/or provide multi-architecture builds.
Finally, benchmarking methods could benefit from high performance computing-style user quotas and timeouts to cap the maximum memory and/or CPU cycles each method can use.
@@
Even a simple, single-person benchmark comparing methods relies on multiple contributions, from dataset creators and method developers to system software and workflow authors. Hence, data, code and operating system dependencies are all accessible during the benchmarking process. For reproducibility, transparency and to build trust [59], this accessibility and freedom should be extended to others to run (re-run), copy (including fork), distribute (e.g., snapshot), study (e.g., read how someone is implementing a method), change (bug fix, contribute) and improve (extend) benchmarks. Such freedoms are granted (or restricted) by software and content licenses, as are copyright and authorship and attribution [60]. Without explicit licenses, even publicly available code is copyrighted by default, and reuse is restricted (to reproduce, distribute, or create derivative works) [60]. A concise summary of license rights and restrictions, and compatibility across commonly-used open-source licenses can be found in Table 1. Benchmark contributions should be licensed and associated to scholarly names including email and ORCID (as the canonical identifier), hence allowing seamless reuse, transparency and attribution. As a consequence, properly-licensed benchmark components (e.g., a procedure to compute a clustering metric; or a clustering method) could be reused across benchmarks; we envision assisted benchmark design tools to leverage and reuse benchmark components, perhaps guided by AI.
@@
Similarly, even if all benchmark code were free and open source (FOSS) software, running it could be hindered by (non-free) proprietary interpreters (e.g., MATLAB). Benchmarking systems should be explicit in allowing or disallowing proprietary software, and be aware and track EULAs (end user license agreements). This conundrum extends to the operating system layer and to system dependencies (e.g., compilers). In terms of benchmark re-runnability, fully FOSS system dependencies arguably offer the highest reusability, and can be handled in a reproducible manner across operating systems and system architectures (e.g., amd64, arm64) [53].
@@
  • Versioning granularity: a single benchmark could address performance of a given method across its versions when run on the same datasets; independently, this benchmark could be updated so several benchmark versions are released. Benchmarking systems should provide versioning at both component and benchmark levels.
@@ Acknowledgments
Funding
MDR acknowledges funding from the Swiss National Science Foundation (grants 200021_212940 and 310030_204869) as well as support from swissuniversities P5 Phase B funding (project 23-36_14). CS is supported by the Novartis Research Foundation. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Conflict of interest
We report no conflict of interest.
@@
Conceptualization: all. Figures: IM, BC. Original draft preparation: all. Writing, reviewing and editing: all. Supervision: IM, MDR. Funding acquisition: MDR.